# Liu He1, Yizhi Song1, Hejun Huang2, Daniel Aliaga1, Xin Zhou2

### v1

@@ Abstract
Text-to-video generation has been dominated by end-to-end diffusion-based or autoregressive models. On one hand, those novel models provide plausible versatility, but they are criticized for physical correctness, shading and illumination, camera motion, and temporal consistency. On the other hand, film industry relies on manually-edited Computer-Generated Imagery (CGI) using 3D modeling software. Human-directed 3D synthetic videos and animations address the aforementioned shortcomings, but it is extremely tedious and requires tight collaboration between movie makers and 3D rendering experts. In this paper, we introduce an automatic synthetic video generation pipeline based on Vision Large Language Model (VLM) agent collaborations. Given a natural language description of a video, multiple VLM agents auto-direct various processes of the generation pipeline. They cooperate to create Blender scripts which render a video that best aligns with the given description. Based on film making inspiration and augmented with Blender-based movie making knowledge, the Director agent decomposes the input text-based video description into sub-processes. For each sub-process, the Programmer agent produces Python-based Blender scripts based on customized function composing and API calling. Then, the Reviewer agent, augmented with knowledge of video reviewing, character motion coordinates, and intermediate screenshots uses its compositional reasoning ability to provide feedback to the Programmer agent. The Programmer agent iteratively improves the scripts to yield the best overall video outcome. Our generated videos show better quality than commercial video generation models in 5 metrics on video quality and instruction-following performance. Moreover, our framework outperforms other approaches in a comprehensive user study on quality, consistency, and rationality.
Introduction
@@
Figure 1: Kubrick Framework and Generated Video Screenshots. Our video generation pipeline consists of multiple agent collaborations, and iterative refinement of both generated visual results and code libraries. The LLM-Director decomposes the user-provided video description into functional sub-processes. Then, for each sub-process LLM-Programmer makes use of a set of function libraries to compose Python scripts for Blender. Intermediate screenshots and video outputs will be reviewed by VLM-Reviewer for visual quality and prompt-following evaluation. The review will be fed back to LLM-Programmer to iteratively improve the video. All three agents are improved by RAG based on public video tutorials and online documents.
@@
Providing AI-assisted video generation is useful to facilitate drafting video sequences and synthetic animations. The video generation process should support high-level storyboard control of the video and handle the details of providing physically-based motions, proper shading and illumination, smooth camera motion, and temporal consistency. Thus, unlike the traditional tedious process of hand-crafted video editing limited to experts, a broader audience of users can focus on the storyboard and benefit from the AI-assisted methodology to produce high-quality videos of varying length and for a variety of content.
@@
Prior methods can be divided into end-to-end video generation approaches (e.g. SORA (2024), Pika (2023), Gen-2 (2024), and Dream Machine (2024)), and synthetic procedural modeling techniques (e.g., Kubric (2022), SceneX (2024), SceneCraft (2024)). The former generate sequences automatically but struggle with physical-correctness, shading and illumination, camera control, and temporal consistency. The latter can address the aforementioned issues but are usually for static 2D or 3D content, and do not address high-level storyboard control.
@@
Our approach defines a pipeline consisting of three intertwined phases (Fig. 1). The Collaborative Generating Phase feeds a user-provided natural-language-based video description to our LLM-Director Agent. This agent in turn produces instructions which are fed to our LLM-Programmer Agent, eventually yielding Python scripts for Blender resulting in a video containing the directed 2D/3D animation. The Visual Reviewing Phase uses a VLM-Reviewer Agent to provide feedback to the LLM-Programmer about the quality of the produced video sequence. This feedback causes the LLM-Programmer to iteratively alter its Python script output until successful generation. The Library Updating Phase automatically finds and/or updates the agent customized code libraries needed by Blender. All the agents use public tutorials and guidelines in a Retrieval Augmented Generation (RAG) (Lewis et al. 2020) based framework to enhance specific knowledge. With our pipeline we have produced numerous videos, under simple natural-language based control, and produce video quality far superior to three comparison baselines Pika (Pika 2023), Gen-2 (Runaway 2024), and Dream Machine (Luma 2024).
@@
Related Works
Video Creation
@@
One recently popular approach for video creation are large-scale training efforts of generative models on video data. Such end-to-end systems use text, image, and natural language descriptions to produce video clips of up to approximately a minute in length. The underlying engine can be diffusion based including, SORA (Brooks et al. 2024), Stable Video Diffusion (Blattmann et al. 2023), VideoPoet (Kondratyuk et al. 2023), and PYoCo (Ge et al. 2023). It can also use transformers like GenTron (Chen et al. 2024), or use divide-and-conquer strategies for better control as Microcinema (Wang et al. 2024a). Some may re-use existing videos (Peng et al. 2024), or image-pose pairs for character animation (Ma et al. 2024). Others can be queued off audio (Yariv et al. 2024) or can focus on particular video categories such as animating people (Hu 2024) or creating dances (Zeng et al. 2024).
However, these end-to-end generation systems struggle to maintain physically-correct motions and behaviors, proper shading and illumination, good camera control, and temporal consistency. Some systems address a few of these issues, such as Bar-Tal et al. focus on temporal consistency and Yang et al. provide separated object and camera controls.
LLMs and Agents
@@
As inspiration for your work, we build upon the reasoning ability of LLMs and improved abilities provided by collaborating agents. As recently shown, LLMs are unexpectedly good at reasoning (2022; 2023; 2024). Moreover, they can be used to help with acting (Yao et al. 2022), generating programming statements (Chen et al. 2022) (Wang et al. 2022), decision making (Yao et al. 2024), and performing iterative refinement (Madaan et al. 2024) without necessarily requiring fine-tuning or re-training (Shinn et al. 2024).
In addition, the use of several collaborating agents has very recently found success in several applications. For example, several agents, or components, collaborate for better planning (Zhang et al. 2024).
Procedural and Synthetic Generation
@@
Our work also builds upon realistic procedural/synthetic scene generation. Broad applications introduce deep learning to visual content generation  (2024; 2023; 2024; 2023; 2023; 2024a; 2024; 2023; 2022; 2020; 2020; 2024b; 2024; 2024). Raistrick et al. produce static 3D scenes of the natural world. Wang et al. generate 3D assets based on exemplars. Deitke et al. procedurally generate physically-enabled scenes. Kubric (Greff et al. 2022), SceneX (Zhou et al. 2024), and SceneCraft (Hu et al. 2024) use 3D modeling software (e.g., Blender) and LLMs to generate 3D static scenes. With these synthetic approaches, obtaining plausibly correct physical structures and good shading and illumination is much easier.
@@
However, these procedural outputs do not produce videos of the generated content and much less according to a user-guided storyboard. Enabling this additional step requires more complex high-level director control and feedback that are not provided previously. Yamada et al. uses a trial-and-error approach to enable feedback but does not produce videos. Zhuang et al. use a diffusion-based approach to create video sequences with some notion of high-level director control, but are not synthetic content and thus suffers from aforementioned limitations of end-to-end video generation.
@@
Method
@@
Collaborative Generating Phase
In this first phase, a LLM-Director agent collaborates with a LLM-Programmer agent (as well as other phases). Designing and then creating a compelling video requires addressing many details. All-at-once solutions may easily lead to unreasonable results. Similar to the chain-of-thoughts (Wei et al. 2022), we observed that decomposing the task to sub-processes benefits video quality and text-to-video alignment. The decomposition may also contribute to effective visual reviewing and to reducing the number of generate-review-edit cycles.
Taking inspiration from film making, the mise-en-scène (i.e., all aspects of the scene/video generation) is addressed by decomposing the generation process into the following five sub-processes (Kawin 1992):
@@
Visual Reviewing Phase
@@
Dynamic Content Evaluation
Unlike former works which evaluate static 3D scene generation (BlenderGPT 2023; Hu et al. 2024), we evaluate character motions and camera motions. As Fig. 3 shows, for the character motion the LLM-Programmer fixes the camera in a location so as to keep the main character 3D assets in frame and in focus. Since the character may have multiple motions, the trajectory of each motion is evaluated separately. Specifically, the VLM-Reviewer samples the starting and ending coordinates of the character’s armature for each motion, text-based location coordinates, motion type, motion speed, and potentially other metrics. In addition, the VLM-Reviewer is provided with screenshots of key frames. Based on the answers of all the evaluation questions, the VLM-Reviewer will give a overall evaluation of ”Pass” or ”Reject”, together with improvement suggestions for the rejected questions. If $v_{i}$ is rejected at iteration $t$, corresponding feedback will be given to LLM-Programmer for the scripting improvement in iteration $t+1$. For example, if the ending coordinates of a motion are not reasonable, LLM-Programmer will take the feedback to modify the ending point arguments ($\varphi_{i}$) when calling corresponding motion function $f_{i}$. Finally, all scripts will be accumulated for each sub-process and used to render a full-length video.
@@
Library Updating Phase
@@
Retrieval-based Instructive Learning
@@
The latest LLMs are well-known for their coding capability in popular programming languages. However, we make use of Retrieval Augmented Generation (RAG) (Lewis et al. 2020) to improve performance in synthetic video generation and Blender-oriented scripting. To this end, we use Blender API documents to improve the API calling correctness of the LLM-Programmer and reduce the number of feedback loops. In addition, to improve the LLM-Director and VLM-Reviewer we use online video tutorials (typically made by Blender experts) related to the user-supplied video description. For each video, we use the provided video subtitles (e.g., English text). We setup the retrieval database after receiving the user-provided video description. Then, RAG is used during the execution of each agent.
@@
Figure 4: Qualitative Comparisons. We show four sets (A to D) of generated videos in the ascending order of video description complexity (displayed below each set). Our method excels in handling character/object motions (highlighted in blue) and complex camera moves (highlighted in orange). Pika and Gen-2 fail to generate reasonable character motions in all four sets, and camera motion is limited. Moreover, Gen-2 misunderstands camera motion prompts as the main character (C and D). Luma fails to follow the character motion (B), or generates inappropriate character (D). See Supplemental for more results.
Experiments
Dataset and Implementation Details
@@
We create 21 example text-based video descriptions to use as input to our method and to the baselines. For those descriptions, we create a 3D resource dataset, Odyssey. The dataset consists of character and scene object 3D assets, and rigged armatures which can be implemented on all characters. We collect 3D assets (with textures) from Sketchfab (2024) and rigged armatures from Mixamo (2024). For RAG database, we collect the subtitles of 261 Blender videos from public video sharing platform, together with official Blender API documentations (2024). Then, using the Odyssey dataset and our method, we generate 21 example videos.
@@
Our pipeline needs online access to GPT-4V (OpenAI 2023) with RAG knowledge (OpenAI 2024) and a working version of Blender. As a side note, our pipeline does not prescribe the length of generated videos. Rather, our storyboarding process enables variable length videos as well as assembling longer videos. However, as videos get longer more review iterations are likely needed. We found a good compromise to be $N_{max}=15$ in our tested videos.
Evaluation Metrics
@@
We evaluate videos using five metrics covering visual quality, fidelity, and description-following ability. For overall visual quality and fidelity, we implement frame-wise Frechet Inception Distance (FID) (Heusel et al. 2017), Frechet Video Distance (FVD) (Unterthiner et al. 2018), and video quality assessment (VQA) score (Wu et al. 2023) in both aesthetic quality (VQA-A) and technical quality (VQA-T). Since FVD and FID evaluate the similarity between our generated videos to high-quality video references, we use the synthetic movie Sintel (Blender.org 2010) as our reference. Sintel is a manually crafted Blender-generated movie and considered to be of professional quality. We randomly sample 21 video clips from the Sintel movie (4 to 5 seconds each). Then we sample 21 same-length video clips from our generated videos for comparisons. Moreover, we use CLIP Similarity (CLIP-SIM) (Wu et al. 2021) as the description-following metric. It evaluates the semantic similarity between the generated video contents and corresponding text descriptions.
Method FVD $\downarrow$ FID $\downarrow$ VQA-A $\uparrow$ VQA-T $\uparrow$ CLIP-SIM $\uparrow$ Gen-2 3463.46 298.82 51.25 47.78 0.2042 Pika 2885.68 298.21 26.06 24.41 0.2044 Luma 2728.90 261.24 40.97 37.73 0.2068 Ours 2691.44 245.43 48.20 53.10 0.2155
Table 1: Quantitative results. Our method outperforms three commercial video generation models on most of the quality, fidelity, and description-following metrics.
Quantitative Comparisons
@@
We compare our model against three baseline models: Gen-2 (Runaway 2024), Pika (Pika 2023) and Dream Machine (Luma 2024). 21 videos are generated by each model using the same set of text descriptions. For fairness, each video is only generated once without cherry-picking.
Quantitative comparisons are shown in Tab. 1. Our model outperforms all baseline models in four metrics, and achieves marginally second place in VQA-A. It demonstrates that our generated videos are with better visual quality and fidelity (FVD, FID, and VQA-T). Higher CLIP-SIM score also highlights the strength of our proposed system on description following (see Fig. 4).
Qualitative Comparisons

### v2

# Kubrick: Multimodal Agent Collaborations for Synthetic Video Generation

@@ Abstract
Text-to-video generation has been dominated by diffusion-based or autoregressive models. These novel models provide plausible versatility, but are criticized for improper physical motion, shading and illumination, camera motion, and temporal consistency. The film industry relies on manually-edited Computer-Generated Imagery (CGI) using 3D modeling software. Human-directed 3D synthetic videos address these shortcomings, but require tight collaboration between movie makers and 3D rendering experts. We introduce an automatic synthetic video generation pipeline based on Vision Large Language Model (VLM) agent collaborations. Given a language description of a video, multiple VLM agents direct various processes of the generation pipeline. They cooperate to create Blender scripts which render a video following the given description. Augmented with Blender-based movie making knowledge, the Director agent decomposes the text-based video description into sub-processes. For each sub-process, the Programmer agent produces Python-based Blender scripts based on function composing and API calling. The Reviewer agent, with knowledge of video reviewing, character motion coordinates, and intermediate screenshots, provides feedback to the Programmer agent. The Programmer agent iteratively improves scripts to yield the best video outcome. Our generated videos show better quality than commercial video generation models in five metrics on video quality and instruction-following performance. Our framework outperforms other approaches in a user study on quality, consistency, and rationality.
@@
Figure 1: Kubrick Framework for Video Generation. Kubrick consists of multiple agent collaborations, and iterative refinement of visual results. The LLM-Director decomposes the user-provided video description into functional sub-processes. Then, the LLM-Programmer compose Python scripts for Blender. Intermediate screenshots and video outputs will be reviewed by VLM-Reviewer for visual quality and prompt-following evaluation. The review will be fed back to VLM-Reviewer to iteratively improve the video. All three agents will be enhanced by RAG from online domain knowledge.
1 Introduction
@@
Providing AI-assisted video generation is useful to facilitate drafting video sequences and synthetic animations. The video generation process should support high-level storyboard control of the video and handle the details of providing physically-based motions, proper shading and illumination, smooth camera motion, and temporal consistency. Thus, unlike the traditional tedious process of hand-crafted video editing limited to experts, a broader audience of users can focus on the storyboard and benefit from the AI-assisted methodology to produce high-quality videos of varying length and for a variety of content [26, 50, 32].
@@
Prior methods can be divided into end-to-end video generation approaches (e.g. SORA [7], Pika [44], Gen-2 [46], and Dream Machine [35]), and synthetic procedural modeling techniques (e.g., Kubric [16], SceneX [75], SceneCraft [24]). The former generate sequences automatically but struggle with physical-correctness, shading and illumination, camera control, and temporal consistency. The latter can address the aforementioned issues but are usually for static 2D or 3D content, and do not address high-level storyboard control.
@@
Our approach defines a pipeline consisting of three intertwined phases (Fig. 2). The Collaborative Generating Phase feeds a user-provided natural-language-based video description to our LLM-Director Agent. This agent in turn produces instructions which are fed to our LLM-Programmer Agent, eventually yielding Python scripts for Blender resulting in a video containing the directed 2D/3D animation. The Visual Reviewing Phase uses a VLM-Reviewer Agent to provide feedback to the LLM-Programmer about the quality of the produced video sequence. This feedback causes the LLM-Programmer to iteratively alter its Python script output until successful generation. The Library Updating Phase automatically finds and/or updates the agent customized code libraries needed by Blender. All the agents use public tutorials and guidelines in a Retrieval Augmented Generation (RAG) [29] based framework to enhance specific knowledge. With our pipeline we have produced numerous videos, under simple natural-language based control, and produce video quality far superior to seven comparison baselines (e.g., Pika [44], Gen-2 [46], Dream Machine [35], and more – see Sec. 4).
@@
2 Related Works
2.1 Video Creation
@@
With booming of DM-based image synthesis [30, 10, 21, 38, 52, 19, 19, 20, 25, 42], video creation also attract dense efforts. Such end-to-end systems use text, image, and natural language descriptions to produce video clips of up to approximately a minute in length. The underlying engine can be diffusion based including, SORA [7], Stable Video Diffusion [3], VideoPoet [28], and PYoCo [14]. It can also use transformers like GenTron [8], or use divide-and-conquer strategies for better control as Microcinema [59]. Some may re-use existing videos [43], or image-pose pairs for character animation [36, 55, 31, 49]. Others can be queued off audio [69] or can focus on particular video categories such as animating people [23, 51, 55] or creating dances [70]. However, these end-to-end generation systems struggle to maintain physically-correct motions and behaviors, proper shading and illumination, good camera control, and temporal consistency. Some systems address a few of these issues, such as  [1] focus on temporal consistency and  [65] provide separated object and camera controls.
2.2 LLMs and Agents
@@
As inspiration for your work, we build upon the reasoning ability of LLMs and improved abilities provided by collaborating agents. As recently shown, LLMs are unexpectedly good at reasoning with chain of thoughts [61, 54]. Moreover, they can be used to help with acting [67], generating programming statements [9, 58], decision making [68], and performing iterative refinement [37] without necessarily requiring fine-tuning or re-training [47]. In addition, the use of several collaborating agents has very recently found success in applications for various modalities [53, 34, 73]. For example, several agents, or components, collaborate for better planning [71].
2.3 Procedural and Synthetic Generation
@@
Our work also builds upon realistic procedural/synthetic scene generation. [45] produce static 3D scenes of the natural world. [60] generate 3D assets based on exemplars. [12] procedurally generate physically-enabled scenes. Kubric [16], SceneX [75], and SceneCraft [24] use 3D modeling software (e.g., Blender) and LLMs to generate 3D static scenes. With these synthetic approaches, obtaining plausibly correct physical structures and good shading and illumination is much easier [2, 74].
@@
However, these procedural outputs do not produce videos of the generated content and much less according to a user-guided storyboard. Enabling this additional step requires more complex high-level director control and feedback that are not provided previously. [64] uses a trial-and-error approach to enable feedback but does not produce videos. [76, 33] use a diffusion-based approach to create video sequences with some notion of high-level director control, but are not synthetic content and thus suffers from aforementioned limitations of end-to-end video generation.
@@
3 Method
Figure 2: Kubrick Framework. Our video generation pipeline consists of multiple agent collaborations and iterative refinement of both generated visual results and code libraries. The LLM-Director decomposes the user-provided video description into functional sub-processes. Then, for each sub-process LLM-Programmer makes use of a set of function libraries to compose Python scripts for Blender. Intermediate screenshots and video outputs will be reviewed by VLM-Reviewer for visual quality and prompt-following evaluation. The review will be fed back to LLM-Programmer to iteratively improve the video. All three agents are improved by RAG based on public video tutorials and online documents.
@@ 3 Method
3.1 Collaborative Generating Phase
In this first phase, a LLM-Director agent collaborates with a LLM-Programmer agent (as well as other phases). Designing and then creating a compelling video requires addressing many details. All-at-once solutions may easily lead to unreasonable results. Similar to the chain-of-thoughts [61], we observed that decomposing the task to sub-processes benefits video quality and text-to-video alignment. The decomposition may also contribute to effective visual reviewing and to reducing the number of generate-review-edit cycles.
Taking inspiration from film making, the mise-en-scène (i.e., all aspects of the scene/video generation) is addressed by decomposing the generation process into the following five sub-processes [27]:
@@ 3.1.2 LLM Programmer
3.2 Visual Reviewing Phase
@@
3.2.1 Dynamic Content Evaluation
Unlike former works which evaluate static 3D scene generation [4, 24], we evaluate character motions and camera motions. As Fig. 4 shows, for the character motion the LLM-Programmer fixes the camera in a location so as to keep the main character 3D assets in frame and in focus. Since the character may have multiple motions, the trajectory of each motion is evaluated separately. Specifically, the VLM-Reviewer samples the starting and ending coordinates of the character’s armature for each motion, text-based location coordinates, motion type, motion speed, and potentially other metrics. In addition, the VLM-Reviewer is provided with screenshots of key frames. Based on the answers of all the evaluation questions, the VLM-Reviewer will give a overall evaluation of ”Pass” or ”Reject”, together with improvement suggestions for the rejected questions. If $v_{i}$ is rejected at iteration $t$, corresponding feedback will be given to LLM-Programmer for the scripting improvement in iteration $t+1$. For example, if the ending coordinates of a motion are not reasonable, LLM-Programmer will take the feedback to modify the ending point arguments ($\varphi_{i}$) when calling corresponding motion function $f_{i}$. Finally, all scripts will be accumulated for each sub-process and used to render a full-length video.
@@ 3.2.1 Dynamic Content Evaluation
3.3 Library Updating Phase
@@
3.4 Retrieval-based Instructive Learning
@@
The latest LLMs are well-known for their coding capability in popular programming languages. However, we make use of Retrieval Augmented Generation (RAG) [29] to improve performance in synthetic video generation and Blender-oriented scripting. To this end, we use Blender API documents to improve the API calling correctness of the LLM-Programmer and reduce the number of feedback loops. In addition, to improve the LLM-Director and VLM-Reviewer we use online video tutorials (typically made by Blender experts) related to the user-supplied video description. For each video, we use the provided video subtitles (e.g., English text). We setup the retrieval database after receiving the user-provided video description. Then, RAG is used during the execution of each agent.
@@
Figure 5: Qualitative Comparisons. Our method excels in handling character/object motions (highlighted in blue) and complex camera moves (highlighted in orange). Pika, Gen-2, and LTX-Video struggle to generate reasonable character motions. While WAN-2.1, Mochi, and CogVideo-X can synthesize character movement, their outputs are heavily distorted and contain significant artifacts. Additionally, all baseline models exhibit limited camera motion. Gen-2 misinterprets camera motion prompts as the main character (right), while Luma fails to accurately follow character motion. See the appendix for more results.
4 Experiments
4.1 Dataset and Implementation Details
@@
We create 21 example text-based video descriptions to use as input to our method and to the baselines. For those descriptions, we create a 3D resource dataset, Odyssey. The dataset consists of character and scene object 3D assets, and rigged armatures which can be implemented on all characters. We collect 3D assets (with textures) from Sketchfab [48] and rigged armatures from Mixamo [39]. For RAG database, we collect the subtitles of 261 Blender videos from public video sharing platform, together with official Blender API documentations [6]. Then, using the Odyssey dataset and our method, we generate 21 example videos.
@@
Our pipeline needs online access to GPT-4V [40] with RAG knowledge [41] and a working version of Blender. As a side note, our pipeline does not prescribe the length of generated videos. Rather, our storyboarding process enables variable length videos as well as assembling longer videos. However, as videos get longer more review iterations are likely needed. We found a good compromise to be $N_{max}=15$ in our tested videos.
4.2 Evaluation Metrics
@@
We evaluate videos using five metrics covering visual quality, fidelity, and description-following ability. For overall visual quality and fidelity, we implement frame-wise Frechet Inception Distance (FID) [22], Frechet Video Distance (FVD) [56], and video quality assessment (VQA) score [63]. Since FVD and FID evaluate the similarity between our generated videos to high-quality video references, we use the synthetic movie Sintel [5] as our reference. Sintel is a manually crafted Blender-generated movie and considered to be of professional quality. We randomly sample 21 video clips from the Sintel movie (4 to 5 seconds each). Then we sample 21 same-length video clips from our generated videos for comparisons. Moreover, we use CLIP Similarity (CLIP-SIM) [62] as the description-following metric. It evaluates the semantic similarity between the generated videos and text descriptions.
Table 1: Quantitative results. We report visual metrics of generated videos based on same set of input text prompt across 7 baselines. Our framework shows the best overall performance of visual quality (FVD, FID) and prompt following (CLIP-SIM). Our VQA score is also competitive to the close-sourced baseline Luma. Best results are in bold, second-best results are underlined.
Method FVD $\downarrow$ FID $\downarrow$ VQA $\uparrow$ CLIP-SIM $\uparrow$
CogV-X 2783.48 302.71 51.10 0.1994
Wan-2.1 3088.93 307.68 48.14 0.2021
LTX 3174.03 306.70 17.79 0.1986
Mochi 2999.18 310.23 33.33 0.1948
Gen-2 3463.46 298.82 30.59 0.2042
Pika 2885.68 298.21 14.04 0.2044
Luma 2728.90 261.24 40.97 0.2068
Ours 2691.44 245.43 40.31 0.2155
4.3 Quantitative Comparisons
@@
We compare our model against seven baseline models: Gen-2 [46], Pika [44], Luma Dream Machine [35], CogVideo-X [66], Wan-2.1 [57], Mochi [15], and LTX-Video [18]. 21 videos are generated by each model using the same set of text descriptions. For fairness, each video is only generated once without cherry-picking.
Quantitative comparisons are shown in Tab. 1. Our model outperforms all baseline models in three metrics and shows competitive performance on VQA score compared to the second best overall close-sourced model Luma. The latest open-sourced model CogVideo-X only outstands on VQA score but regresses on all other metrics. It demonstrates that our generated videos keep the balance of visual quality &\&& fidelity (FVD and FID), and better prompt following indicated by higher CLIP-SIM score (see Fig. 5).
Table 2: User Study. We compare different text-to-video generation methods in four criteria. Each user will watch 2 random videos generated by each model, and rate the score of each criteria in the scale from 1 to 5. Below shows the average score of each criteria. Our framework shows the best overall performance on all criteria by a significant margin. Best results are in bold, second-best results are underlined.
Method Evaluation Criteria
Prompt $\uparrow$ Motion $\uparrow$ Camera $\uparrow$ Overall $\uparrow$
CogV-X 2.95 2.43 2.47 2.52
Wan-2.1 3.48 3.28 3.30 3.35
LTX 2.27 1.85 2.22 1.87
Mochi 3.47 2.88 3.07 3.13
Pika 2.97 2.65 2.85 2.87
Gen-2 2.67 2.38 2.60 2.52
Luma 3.47 3.28 3.35 3.32
Ours 4.18 3.83 4.02 3.93
4.4 Qualitative Comparisons
Evaluation Visualization